OpenStudy (anonymous):

Flourishing Flowers charges a flat fee of $125 plus $60 for each standard floral arrangement to deliver and set up flowers for a banquet. Find the cost of providing 20 floral arrangements. a. $875 c. $1325 b. $950 d. $1275

OpenStudy (anonymous):

????

OpenStudy (anonymous):

anyone?

jimthompson5910 (jim_thompson5910):

Total Cost = Flat Fee + (Variable Fee) * (Number of Items) Total Cost = 125 + 60*20 Total Cost = ???

OpenStudy (anonymous):

3700

jimthompson5910 (jim_thompson5910):

do you have a calculator?

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

ok, you multiply BEFORE you add

jimthompson5910 (jim_thompson5910):

sry the order is a bit confusing i know, but PEMDAS says that multiplication comes first, then addition

OpenStudy (anonymous):

120 + 125 = 245

jimthompson5910 (jim_thompson5910):

Total Cost = 125 + 60*20 Total Cost = 125 + 1200 Total Cost = ???

OpenStudy (anonymous):

oh du 1200 okay so 1325

jimthompson5910 (jim_thompson5910):

you got it